I have no clue how to solve Systems of Equations!!!??? HELP EX: y=3x+6
y=x+2
Any Suggetions??

Answer by ankor@dixie-net.com(22740)

y = 3x + 6
y = x + 2
this one is easy to solve using substitution
we know that y = y, therefore we can write it:
3x + 6 = x + 2
3x - x = 2 - 6
2x = -4
x = -4/2
x = -2
:
Find y
y = 3x + 6
replace x with 2
y = 3(-2) + 6
y = -6 + 6
y = 0
:
Check solutions in the 2nd equation
y = x + 2
Replace x and y and see if it is equal
0 = -2 + 2